Janny opened a savings account and deposited $300.00 as principal. The account earns 6% interest, compounded annually. What is the balance after 6 years?

1
Expert's answer
2021-12-20T16:32:41-0500

By compound interest:

"F=P(1+r)^n=300(1+0.06)^6=425.56"